OpenStudy (anonymous):

Emma has a rope that is square root sign-180 cm long. She wants to cut the rope into square root sign-45 cm long pieces. How many pieces can Emma make?

OpenStudy (mertsj):

\[\frac{\sqrt{180}}{\sqrt{45}}=\sqrt{\frac{180}{45}}=\sqrt{4}=2\]